OpenStudy (anonymous):

HI EVERYONE!! Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? A) x + 2 = -4 B) -x - 2 = 4 C) x + 2 = 4 D) x - 2 = -4

OpenStudy (michele_laino):

please you have to consider two distinct cases, namely: \[x-2\ge 0, x-2=4\] and: \[x-2<0,-x+2=4\]

OpenStudy (anonymous):

what is your choice?

OpenStudy (anonymous):

just open \[\left| x-2 \right|\]

OpenStudy (michele_laino):

please I rewrite: case 1) \[x-2\ge 0\] then your equation can be rewritten as below: \[x-2=4\] case 2) \[x-2< 0\] then your equation can be rewritten as below: \[-x+2=4\]

OpenStudy (anonymous):

exactly

OpenStudy (michele_laino):

@M_A thanks!

OpenStudy (anonymous):

@kelliegirl33 @Michele_Laino @M_A Thank you all for your time and efforts.

OpenStudy (anonymous):

its either : (x - 2) = 4 OR -(x - 2) = 4

OpenStudy (michele_laino):

@extremearianna thank you!

OpenStudy (anonymous):

@kelliegirl33 what happened to the other comment? I thought it might have been D

OpenStudy (michele_laino):

your answer is D

OpenStudy (anonymous):

Thank you very much! I will fan you also :) @Michele_Laino

OpenStudy (michele_laino):

Thank you again!